OpenStudy (anonymous):

During a hike, Vijay saw a total of 15 birds and butterflies. Knowing that birds have 2 wings and butterflies have 4 wings, he counted 48 wings. Use a model to determine how many birds he saw. (3 points) During a hike, Vijay saw a total of 15 birds and butterflies. Knowing that birds have 2 wings and butterflies have 4 wings, he counted 48 wings. Use a model to determine how many birds he saw. (3 points) @Mathematics

OpenStudy (anonymous):

6

OpenStudy (anonymous):

may be 12 birds and 3 butterflies

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

No 12 and 3 doesn't seem right

OpenStudy (anonymous):

can u tell me which is the rit ans coz i thought 24 wings means 12 birds and 12 wings means 3birds

OpenStudy (anonymous):

Wings should total to 48

OpenStudy (anonymous):

when u will divide 48 wing by 3 u get 12 wingss and wen u divide it by 2 u get 24 wings

OpenStudy (anonymous):

the answer should b 6 right?

OpenStudy (anonymous):

birds = 6 and butterflies = 9

OpenStudy (anonymous):

24 wings means 12 birds and 12 wings means 3butterfly

OpenStudy (anonymous):

na its 6 i just did the math heena

OpenStudy (anonymous):

i m also not sure i m also just doing hit and try but can u tell me how it is 6 and 9

OpenStudy (anonymous):

no. of birds = x no. of butterflies = y x + y =15 y = 15-x 2x + 4y = 48 2x+ 4(15-x) =48 2x + 60-4x = 48 -2x +60 = 48 60 - 48 = 2x 12 = 2x x = 6 y = 9

OpenStudy (anonymous):

yes like that

OpenStudy (anonymous):

ohkeezz.. :-) i have left math since 2 year so i totaly forgot about quadratic equation

OpenStudy (anonymous):

ok :)